# Expanding Deep Learning-based Sensing Systems with Multi-Source Knowledge Transfer

### v1

@@ 4.1. Efficient Model Selection Protocol
Lightweight features $l=\{l^{1},l^{2},...\}$, including the mean value, standard deviation, skewness, and kurtosis, are extracted from both the source domains and target domain data. These features provide a coarse-grained description of domain characteristics without relying on any learning models. The target domain transmits its extracted features to the source domains, where each source domain $i$ computes a feature similarity score: $s_{i}=1/J\sum_{j=1}^{J}sim(l_{i}^{j},l_{t}^{j})$ where $J$ is the total number of features. The similarity function $sim(\cdot)$ is set to cosine similarity. The target domain then selects the top $\eta\%$ of source domains with the highest similarity scores. Only the selected domains that receive a model inquiry from the target domain transmit their models back. This approach significantly reduces communication traffic, as only a small percentage of models are transmitted and high-level features are much smaller in size compared with those models.
Figure 4. The Efficient Model Selections Protocol.
4.1.2. Centroids-Accuracy Joint Selection.
@@
After the models are received in the target domain, they are further selected based on their accuracy and the similarity of their class centroids. The accuracy $Acc_{i}$ of the $i$-th source model is computed on the target domain labeled data, which, however, can be unreliable due to the limited amount of labeled data. To address this, we further use class centroids, which represent the class-wise feature distributions captured by the models. These centroids are computed by averaging the encoder outputs for samples within each class:
(1) $\begin{split}c_{t,m}=\frac{1}{K}\sum_{k=1}^{K}{h_{i}(k)},k\in\{k|max(g_{i}(h_{i}(k)))=m\\ \text{and }rank(e(k))<75\%\}\end{split}$
@@
where $h_{i}(k)=f_{i}(x(k))$ represents the features extracted by the encoder $f_{i}$ of the $i$-th domain from the $k$-th data sample. The classifier output, $g_{i}(h_{i}(k))$, is processed using the function $max(\cdot)$ to obtain a pseudo label. However, due to data heterogeneity, the pseudo labels produced by $g_{i}$ and $f_{i}$ on the target domain may be inaccurate. To enhance the quality of the centroids, entropy, a measure of the prediction confidence, is used to filter out features with low certainty. The entropy $e(k)$ of the logits output $g_{i}(f_{i}(k))$ is calculated, and the $75\%$ of features with the lowest entropy (i.e., the highest confidence) are selected for centroid extraction.
The centroids of the source and target domains are compared to estimate the data distribution similarity $s_{fine}(t,i)$:
(2) $s_{fine}(t,i)=\frac{1}{M}\sum_{m=1}^{M}\text{sim}(c_{t,m},c_{i,m}),$
@@
where $M$ is the number of overlapping classes between the source and target domains. A higher value of $s_{fine}(t,i)$ indicates a greater similarity in data distribution between the $i$-th source and target domains, suggesting that the source model is likely to experience a smaller performance drop when applied to the target data. Compared to selection methods (Li et al., 2019; Borup et al., 2023) that rely solely on labeled data, using centroids extracted from both labeled and unlabeled data provides a more accurate reflection of domain similarity and model effectiveness. The centroid similarity score $s_{fine}(t,i)$ is then multiplied by the model accuracy $Acc_{i}$ to rank the source domains and determine the final set of $N_{p}$ selected models.
4.2. Sample-wise Knowledge Fusion
@@
To resolve knowledge conflicts, an attention-based mixer is trained to assign sample-wise weights to the selected models based on their relative importance. Simultaneously, a cost-effective adaptation is applied to the source models to enhance their knowledge quality.
4.2.1. Attention-based Mixer.
@@
The attention-based mixer aggregates conflicting predictions from the selected models by leveraging the sample-wise feature adjacency between the source and target models. The feature $h_{t}(k)$ extracted by the target encoder is projected through a linear layer $L^{\text{query}}$ to obtain the query vector $q(k)$. Similarly, the features $h_{i}(k)$, extracted by the selected source models, are projected through the respective linear layers $L_{i}^{\text{key}}$ to obtain the key vectors $key_{i}(k)$:
(3) $q(k)$ $=L^{\text{query}}(h_{t}(k)),$
(4) $key_{i}(k)$ $=L_{i}^{\text{key}}(h_{i}(k)).$
@@
Different from the traditional attention mechanism (Vaswani, 2017), which uses a single linear layer to compute the keys, the mixer utilizes multiple linear layers $L_{i}^{\text{key}}$ to accommodate the heterogeneous source model architectures. Since the models from different source domains extract features with varying dimensions, the input size of each $L_{i}^{\text{key}}$ must be tailored accordingly. The output size of $L_{i}^{\text{key}}$ is standardized to a common dimension for subsequent computations. The similarities between the query and the keys are then calculated and normalized using SoftMax to obtain the attention score $w_{i}(k)$ for the $i$-th model on the data sample $x(k)$:
(5) $w_{i}(k)=\text{SoftMax}(q(k)\cdot key_{i}(k)),\quad i=1,2,\ldots,N_{p}$
@@
(6) $p^{\text{mix}}(k)=\sum_{i}w_{i}(k)g_{i}(h_{i}(k)).$
If the features extracted by the target model and the $i$-th model are highly similar, a higher weight $w_{i}(k)$ is assigned to the prediction of the $i$-th model. This is because the $i$-th classifier is likely to be more accurate on data from a distribution that closely resembles the data it is trained with. However, training the attention-based mixer is challenging, as it depends on input from the target model, which itself requires training. To address this, a low-cost joint training scheme is proposed to train the mixer and the target model simultaneously, which is detailed in Section 4.3.
Figure 5. Details of the Sample-wise Knowledge Fusion and the Adaptive Knowledge Injection process.
4.2.2. Cost-effective Adaptation
To further improve the accuracy of the fused predictions, a cost-effective adaptation is applied to the selected models, enhancing the prediction accuracy of each. Since adapting all $N_{p}$ selected source models would be computationally expensive, only their classifiers are trained jointly with the mixer, while their encoders remain frozen. This approach reduces the computational burden, as classifiers are typically lightweight (He et al., 2016). Early-stage experiments indicate that adapting the classifiers alone is sufficient to provide high-quality predictions for the attention-based mixer. Additionally, freezing the encoders accelerates the knowledge aggregation process. By precomputing features for all target domain data using the frozen encoders and storing them in memory, they are ready to be fetched when the mixer requires the prediction results for a sample $x(k)$. This process eliminates the need to repeatedly execute the forward pass of the selected encoders, thereby reducing the overall computation time.
4.3. Adaptive Knowledge Injection
The fused knowledge is further distilled into the target model:
(7) $L_{\text{ada}}=L^{\text{label}}+\alpha\,L^{\text{distill}}(g_{t}\cdot f_{t}(x(k)),KD(k)),$
@@ 4.3. Adaptive Knowledge Injection
4.3.1. Knowledge Dictionary.
Directly learning from the fused prediction $p^{\text{mix}}$ can hinder the convergence of the target model. This is because the aggregation results from the mixer change dynamically during its training (e.g., the pseudo-label of a sample may shift from class A to class C), potentially leading to conflicting gradient update directions. Additionally, the quality of the fused predictions may fluctuate from epoch to epoch. To address this issue, a knowledge dictionary is introduced to provide more stable learning objectives. After each model update, the fused predictions from the attention-based mixer are stored in the knowledge dictionary only if the accuracy of the mixer improves. These predictions are stored in a soft-label format rather than as one-hot vectors, allowing the target model to capture the confidence levels of the mixer in its fused predictions. In subsequent epochs, the target model learns from these soft pseudo labels stored in the knowledge dictionary, rather than directly from the potentially unstable predictions generated by the mixer.
Algorithm 1 Low-cost Joint Training of HaKT
1:  Input: Target labeled and unlabeled data $X^{\{l,u\}}$, labels $Y^{l}$, selected encoders and classifiers $\{f_{i},g_{i}\},$
@@
3:  Initial $g_{t},f_{t},Mixer,KD$
4:  $\{h_{i}^{\{l,u\}}\}$ = Encoding($X^{\{l,u\}}$, $\{f_{i}\}$)
@@
7:     $Mixer$ = Mixer_Update($\{h_{i}^{l}\},h_{t}^{l},Mixer,Y^{l}$)
@@
4.3.2. Adaptive Learner.
Given the varying quality of the fused predictions in the knowledge dictionary, an adaptive learner is employed to adjust the weight $\alpha$ of the distillation loss: $\alpha=m(Acc_{\text{train}}-b)$, where $Acc_{\text{train}}$ represents the accuracy of the attention-based mixer on the training data. The $m$ and $b$ are predetermined hyperparameters. The $m$ controls the scaling factor of the weight $\alpha$, while $b$ serves as a threshold to prevent the target model from learning from fused predictions of low quality. The weight $\alpha$ increases when the fused prediction accuracy is high, allowing the model to learn more effectively from reliable predictions.
4.3.3. Low-cost Joint Training
The training of the attention-based mixer, the unfrozen source classifiers, and the target domain model is complex, as they depend on inputs from each other. To enable a cost-effective training process, a joint training scheme is developed as shown in Algorithm 1.
The labeled and unlabeled data, $X^{\{l,u\}}$, are encoded by the frozen source encoders $\{f_{i}\}$ to high-level features, which are later used to train both the attention-based mixer and the target model. In each epoch, the target encoder $f_{t}$ processes the labeled and unlabeled data, generating feature representations $h_{t}^{l}$ and $h_{t}^{u}$. To manage the computational overhead of encoding a large volume of unlabeled data $X^{u}$ and assigning pseudo labels, only a subset of $X^{u}$ is sampled in each epoch, with the sample size kept proportional to the size of the labeled data. This strategy ensures that the entire set of unlabeled data is progressively utilized over multiple iterations, thereby reducing training time and memory usage of each epoch without compromising model performance. The knowledge dictionary is updated only when the mixer’s quality improves, minimizing the cost of generating pseudo-labels for all unlabeled data.
@@ 4.3.3. Low-cost Joint Training
5. Evaluations
5.1. Experiment Setting
5.1.1. Datasets
Table 1. Datasets for evaluating the effectiveness and generalizability of HaKT.
Dataset Class Domain (#) Task Modality Model Library
Audio, Radar
HaKT is evaluated on four datasets that span various modalities, tasks, and scales. Table 1 provides a summary of these datasets, with further details as follows:
Alzheimer’s Disease (AD) (Ouyang et al., 2023). This dataset consists of Alzheimer’s Disease-related activity data collected from 16 home environments using multiple sensing modalities. It includes 11 activity classes, such as writing and sleeping.
While HaKT is evaluated on these three diverse applications, it has the potential to extend to other sensing systems, such as traffic management or smart agriculture (Jiang et al., 2023; Zhu et al., 2018), which we plan to explore in future work.
5.1.2. Model Libraries.
Table 1 summarizes the model libraries, which include six different models for each dataset. For IMU data, the TPN-(S, M, L) (Saeed et al., 2019) and CPC-(S, M, L) (Haresamudram et al., 2021) models are used, with feature channels of 12, 16, 32 for TPN and 8, 12, 16 for CPC, respectively. For image processing, the model library consists of GoogleNet (Szegedy et al., 2015), MobileNet-v3 (S, L) (Howard et al., 2019), and ResNet-(18, 34, 50) (He et al., 2016). For EMG data, the ConvNet-(S, M, L) (Côté-Allard et al., 2019) models and a RNN-(S, M, L) are utilized, with feature channels of 4, 8, 12 for ConvNet and 32, 48, 56 for RNN, respectively. To handle the multi-modal data in AD, we adapt the model in (Ouyang et al., 2023) by varying the number of layers and feature dimensions, creating 5-layer ADNet-(S, M, L) and 3-layer TinyADNet-(S, M, L) models with 64, 128, 256 for ADNet and 32, 64, 96 feature channels for TinyADNet.
5.1.3. Baselines
The five most relevant baselines are selected and slightly adapted for comparison:
AccDistill (Li et al., 2019). Source domain models with top-k accuracy are selected and ensembled in (Li et al., 2019). To support model customization, we modify it by transferring the knowledge from the ensembled model to the target models.
MEHLSoup (Li et al., 2024). MEHLSoup merges multiple source domain models with a learned mixing coefficient, which is optimized by a block coordinate gradient descent algorithm on the target domain data.
Figure 6. Performance comparison with different target models. The $\gamma$ is fixed to 10%.
Figure 7. Performance comparison with varied portions of labeled data in the target domain.
Other knowledge distillation, domain adaptation, or model merging methods are not included, as they have already been outperformed by the considered baselines (Borup et al., 2023; Qu et al., 2024; Li et al., 2024). Since LEAD and MEHLSoup, as well as other adaptation and merging methods, are unable to handle model heterogeneity, they are not directly comparable to HaKT. To make both methods executable, we select source domains with architectures that match the target models as candidates. Federated learning methods are not included for comparison due to the difference in the considered scenario (See Section 2.1).
5.1.4. Real-world Testbed.
@@ 5.1.3. Baselines
The system is deployed on a server equipped with a 12th Gen Intel(R) Core i9-12900KF processor and an edge device, Nvidia Jetson Xavier. To simulate different source domains, the source domain data and models are stored in separate folders on the server due to the limited number of available devices. The target domain data is deployed on the edge device. The model training overhead, including time and memory usage, is measured on the edge devices, which are closely correlated with energy consumption, particularly on edge devices. The communication overhead is monitored by tracking the network traffic between the server and the edge device. The real-life deployment on other devices, such as smartphones or personal computers, is further discussed in Section 6.
5.1.5. Implementation Details.
[…]
5.2. Result Comparisons in One-Time Sensing Systems Expansion
5.2.1. Data Split and Training Details.
We first compare the results of HaKT against the baselines in a one-time expansion setting, where one domain is randomly selected as the target domain, and the remaining domains serve as source domains. The source domain architectures are randomly selected from TPN-(S, M, L), ResNet-(18, 34, 50), ConvNet-(S, M, L), and ADNet-(S, M, L). The source domain models are trained using supervised learning on the labeled data of each domain. For the target domains, 60% of the data is randomly selected as the training set, 20% as the validation set, and the remaining 20% as the test set. Only $\gamma\%$ of the training set data is labeled, while the rest remains unlabeled. For ImageNet-R, each of the eight styles is tested separately. For the other three datasets, ten different splits are randomly generated, and the average accuracy and communication overhead are reported.
5.2.2. Vaired Target Architectures
The results on different target architectures in the model library are presented in Figure 6. Notably, LEAD and MEHLSoup can only leverage source domains that share the same architecture as the target domains, making them inapplicable when the required architecture (e.g., CPC-S for HARBox) is not available among the source models. HaKT consistently achieves comparable or superior accuracy across four datasets for most architectures. Specifically, HaKT surpasses the best baseline by 7.0%, 16.5%, 15.9%, 6.4%, 5.0%, and 4.0% for CPC-(S, M, L) and TPN-(S, M, L), respectively. On ImageNet-R, when the target models are MobileNet-(S, L), HaKT achieves gains of 4.1% and 3.4% over the best baselines, respectively. Although HaKT performs slightly worse than MEHLSoup on NinaPro, it significantly outperforms MEHLSoup on the other datasets, likely due to the lower data heterogeneity in NinaPro, which aligns better with MEHLSoup’s approach. Unlike domain adaptation and model merging methods that are limited to source models matching the target architecture, HaKT can effectively utilize knowledge from source models with diverse architectures, leading to improved performance. These results demonstrate HaKT’s versatility in delivering high-quality customized models across diverse tasks and modalities.
@@ 5.2.2. Vaired Target Architectures
Table 2. Performance comparison in the multi-round expansion setting. Accuracy refers to the average accuracy of four rounds of expansion. Traffic is the total communication cost during expansion.

### v2

@@ 4.2 Efficient Model Selection Protocol
Feature-based Coarse Selection. Lightweight features $l=\{l(n)\}$, such as mean value, standard deviation, skewness, and kurtosis, are extracted from both the source and target domain data. These features provide a coarse-grained description of domain characteristics without relying on any learning models. The target domain transmits its extracted features to the source domains, where each source domain $i$ computes a feature similarity score: $s_{i}=\sum_{n=1}^{N}sim(l^{S}_{i}(n),l_{t}(n))/N,$ where $N$ is the total number of features. The similarity function $sim(\cdot)$ is set to cosine similarity. The target domain then selects the top $\eta\%$ of source domains with the highest similarity scores. Only the selected domains that receive a model inquiry from the target domain transmit their models back.
@@
Centroids-Accuracy Joint Selection. After the models are received in the target domain, the accuracy $Acc_{i}$ of the $i$-th source model is computed on the target domain labeled data, which, however, can be unreliable due to label scarcity. To address this, we further use class centroids, which represent the class-wise feature distributions captured by the models. These centroids are computed by averaging the encoder outputs for samples within each class:
$\begin{split}c_{t,m}=\frac{1}{K}\sum_{k=1}^{K}{h_{i}(k)},k\in\{k|max(g^{S}_{i}(h_{i}(k)))=m\\ \text{and }rank(e(k))<\omega\}\end{split}$ (1)
@@
where $h_{i}(k)=f^{S}_{i}(x(k))$ represents the features extracted by the encoder $f^{S}_{i}$ of the $i$-th domain from the $k$-th data sample. The classifier output, $g^{S}_{i}(h_{i}(k))$, is processed with $max(\cdot)$ to obtain a pseudo label, which may be inaccurate due to data heterogeneity. To enhance the quality of the centroids, entropy, a measure of the prediction confidence, is used to filter out features with low certainty. The entropy $e(k)$ of the logits output $g^{S}_{i}(h_{i}(k))$ is calculated, and the $\omega=75\%$ of features with the lowest entropy (i.e., the highest confidence) are selected for centroid extraction.
@@
The centroids of $\mathcal{D}_{t}$ and $\mathcal{D}^{S}_{i}$ are compared to estimate the data distribution similarity $s^{fine}(t,i)$: $s^{fine}(t,i)=\sum_{m=1}^{M}\text{sim}(c_{t,m},c^{S}_{i,m})/M,$ where $M$ is the number of overlapping classes between $\mathcal{D}_{t}$ and $\mathcal{D}^{S}_{i}$. A higher value of $s^{fine}(t,i)$ indicates a greater similarity in data distribution, suggesting that the source model is likely to experience a smaller performance drop when applied to the target data. The centroid similarity score $s^{fine}(t,i)$ is then multiplied by the accuracy $Acc_{i}$ to rank the source domains and determine the final set of $N_{p}$ selected models.
This approach reduces communication traffic and inference costs, as only a subset of models is transmitted and tested in $\mathcal{D}_{t}$, given that features are much smaller in size compared to models. Compared to methods that require access to each sample in $\mathcal{D}_{t}$ and $\mathcal{D}^{S}_{i}$ (He et al., 2023; Du et al., 2024), HaT enhances privacy by exchanging only high-level features and models between domains. Furthermore, privacy can be further strengthened by selectively sharing non-sensitive features while excluding sensitive content (Qu et al., 2024).
4.3 Sample-wise Knowledge Fusion
@@
To resolve knowledge conflicts, an attention-based mixer is trained to assign sample-wise weights. Simultaneously, a cost-effective adaptation is applied to the source models to enhance their knowledge quality.
@@
Attention-based Mixer. The attention-based mixer aggregates conflicting predictions from the selected models by leveraging the sample-wise feature adjacency between the source and target models. The feature $h_{t}(k)$ extracted by the target encoder is projected through a linear layer $L^{\text{query}}$ to obtain the query vector $q(k)$. Similarly, the features $h_{i}(k)$, extracted by the selected source models, are projected through the respective linear layers $L_{i}^{\text{key}}$ to obtain the key vectors $key_{i}(k)$:
$q(k)$ $=L^{\text{query}}(h_{t}(k)),$ (2)
$key_{i}(k)$ $=L_{i}^{\text{key}}(h_{i}(k)).$ (3)
@@
Different from the classic attention mechanism (Vaswani, 2017), which uses a single linear layer to compute the keys, the mixer utilizes multiple linear layers $L_{i}^{\text{key}}$ to accommodate the heterogeneous source model architectures. Since models from different source domains extract features of varying dimensions, the input size of each $L_{i}^{\text{key}}$ must be tailored accordingly. The output size of $L_{i}^{\text{key}}$ is standardized to a common dimension for subsequent computations. The similarities between the query and the keys are then calculated and normalized using SoftMax to obtain the attention score $w_{i}(k)$ for the $i$-th model on the data sample $x(k)$:
$w_{i}(k)=\text{SoftMax}(q(k)\cdot key_{i}(k)),\quad i=1,2,\ldots,N_{p}$ (4)
@@
$p^{\text{mix}}(k)=\sum_{i}w_{i}(k)\cdot Map[g^{S}_{i}(h_{i}(k))],$ (5)
where $Map[\cdot]$ projects the label spaces of different domains to a unified label space that includes all categories. If the features extracted by the target model and the $i$-th model are highly similar, a higher weight $w_{i}(k)$ is assigned to the prediction of the $i$-th model. This is because the $i$-th classifier is likely to be more accurate on data from a distribution that closely resembles the data it is trained with. However, training the attention-based mixer is challenging, as it depends on input from the target model, which itself requires training. To address this, a low-cost joint training scheme is proposed to train the mixer and the target model simultaneously, which is detailed in Section 4.4.
Figure 4: Details of the Sample-wise Knowledge Fusion and the Adaptive Knowledge Injection process.
Cost-effective Adaptation To further improve the accuracy of the fused predictions, the classifiers are trained jointly with the mixer, while their encoders remain frozen, since adapting all $N_{p}$ selected source models would be computationally expensive. This approach reduces the computational burden, as classifiers are typically lightweight (He et al., 2016). Early-stage experiments indicate that adapting the classifiers alone is sufficient to provide high-quality predictions for the attention-based mixer. Additionally, freezing the encoders accelerates the knowledge aggregation process. By precomputing features for all target domain data using the frozen encoders and storing them in memory, they are ready to be fetched when the mixer requires the prediction results for a sample $x(k)$. This process eliminates the need to repeatedly execute the forward pass of the selected encoders, thereby reducing the overall computation time.
4.4 Adaptive Knowledge Injection
Considering the constraints $\zeta_{t}$ on memory and the inference speed of the target, HaT selects a suitable model skeleton from a model library through device profiling. The model library includes open-source models from the Internet, the architectures from $\mathcal{D}^{S}$, or models contributed by the target. Then, the fused knowledge is distilled into the target model:
$L_{\text{ada}}=L^{\text{label}}+\alpha\,L^{\text{distill}}(g_{t}\cdot f_{t}(x(k)),KD(k)),$ (6)
@@ 4.4 Adaptive Knowledge Injection
Knowledge Dictionary. Directly learning from the fused prediction $p^{\text{mix}}$ may hinder the convergence of the target model. This is because the aggregation results from the mixer change dynamically during its training (e.g., the pseudo-label of a sample may shift from class A to C), leading to conflicting gradient update directions. Additionally, the quality of the fused predictions may fluctuate from epoch to epoch. The knowledge dictionary $KD$ is introduced to provide more stable learning objectives. After each model update, the fused predictions from the attention-based mixer are stored in the $KD$ only if the accuracy of the mixer improves. These predictions are stored in a soft-label format rather than as one-hot vectors, allowing the target model to capture the confidence levels of the mixer in its fused predictions. In subsequent epochs, the target model learns from the soft pseudo labels in the $KD$.
Adaptive Learner. Given the varying quality of the fused predictions in the knowledge dictionary, an adaptive learner is employed to adjust the weight $\alpha$ of the distillation loss: $\alpha=m(Acc_{\text{train}}-b)$, where $Acc_{\text{train}}$ represents the accuracy of the attention-based mixer on the training data. The $m$ and $b$ are predetermined hyperparameters. The $m$ controls the scaling factor of the weight $\alpha$, while $b$ serves as a threshold to prevent the target model from learning from fused predictions of low quality. The weight $\alpha$ increases when the fused prediction accuracy is high, allowing the model to learn more effectively from reliable predictions.
Low-cost Joint Training The training of the attention-based mixer, the unfrozen source classifiers, and the target domain model is complex, as they depend on inputs from each other. To enable a cost-effective training process, a joint training scheme is developed as shown in Algorithm 1.
Algorithm 1 Low-cost Joint Training of HaT
1:  Input: Target labeled and unlabeled data $X^{\{l,u\}}$, labels $Y^{l}$, selected models $\{f^{S}_{i}\}_{i=1}^{N_{p}},\{g^{S}_{i}\}_{i=1}^{N_{p}},$
@@
3:  Initialize $g_{t},f_{t},Mixer,KD$
4:  $\{h_{i}^{\{l,u\}}\}$ = Encoding($X^{\{l,u\}}$, $\{f^{S}_{i}\}_{i=1}^{N_{p}}$)
@@
7:     $Mixer$ = Mixer_Update($\{h_{i}^{l}\},h_{t}^{l},Mixer,Y^{l},\{g^{S}_{i}\}$)
@@
The labeled and unlabeled data, $X^{\{l,u\}}$, are encoded by the frozen source encoders $\{f^{S}_{i}\}_{i=1}^{N_{p}}$ to high-level features, which are kept for later training. In each epoch, the $f_{t}$ encodes $X^{\{l,u\}}$, generating representations $h_{t}^{l}$ and $h_{t}^{u}$. To manage the computational overhead of processing a large volume of unlabeled data $X^{u}$, only a subset of $X^{u}$ is randomly sampled in each epoch, with the sample size kept proportional to the size of the labeled data. This strategy ensures that the entire set of unlabeled data is progressively utilized over multiple iterations, thereby reducing training time and memory usage of each epoch without compromising model performance. The $KD$ is updated only when the mixer’s quality improves, minimizing the cost of generating pseudo-labels for all unlabeled data.
@@
5 Evaluations
5.1 Experiment Setting
Datasets HaT is evaluated on four datasets, HARBox(Ouyang et al., 2021), ImageNet-R(Hendrycks et al., 2021), NinaPro(Pizzolato et al., 2017), and Alzheimer’s Disease (AD)(Ouyang et al., 2023) that span six modalities, three tasks, and different scales. For each dataset, six different model architectures are included as model libraries, with further details introduced in Appendix A.1 and A.2.
Baselines The five most relevant baselines from knowledge distillation, model aggregation, and domain adaptation are implemented and slightly adapted for comparison, including DistillWeighted (Borup et al., 2023), DistillNearest (Borup et al., 2023), LEAD (Qu et al., 2024), and MEHLSoup (Li et al., 2024a). We also propose a baseline called AccDistill, which distills the knowledge from ensemble models based on selected models from the source domains. The details of baselines are further introduced in Appendix A.3.
@@ 5.1 Experiment Setting
Real-world Testbed. The system is deployed on a server and an edge device, the Nvidia Jetson Xavier. To simulate different source domains, the data and models are stored in separate folders on the server due to the limited number of available devices. The target domain data is deployed on the edge device. The model training overhead, including time and memory usage, is measured on the edge devices, which are closely correlated with energy consumption. The communication overhead is monitored by tracking the network traffic between the server and the edge device.
@@
Table 1: Performance comparison in the MRSE setting. Accuracy refers to the average accuracy of four rounds of expansion. Traffic is the total communication cost during expansion.